# Decoding Geometric Properties in Non-Random Data from First Information-Theoretic Principles

### v2

@@ Abstract
Based on the principles of information theory, measure theory, and theoretical computer science, we introduce a univariate signal deconvolution method with a wide range of applications to coding theory, particularly in zero-knowledge one-way communication channels, such as in deciphering messages from unknown generating sources about which no prior knowledge is available and to which no return message can be sent. Our multidimensional space reconstruction method from an arbitrary received signal is proven to be agnostic vis-a-vis the encoding-decoding scheme, computation model, programming language, formal theory, the computable (or semi-computable) method of approximation to algorithmic complexity, and any arbitrarily chosen (computable) probability measure of the events. The method derives from the principles of an approach to Artificial General Intelligence capable of building a general-purpose model of models independent of any arbitrarily assumed prior probability distribution. We argue that this optimal and universal method of decoding non-random data has applications to signal processing, causal deconvolution, topological and geometric properties encoding, cryptography, and bio- and technosignature detection.
keywords:
algorithmic information dynamics, causal deconvolution, zero-knowledge communication, technosignatures, signal processing, perturbation analysis, universal distribution, intelligent signal detection, biosignatures
1 Introduction
@@
In this case in which the message is encoded in an uncompressed form, or equivalently in which the message is compressible (therefore non-random), one would have a large standard deviation in the range of information values for the different ways of decomposing, parsing or perturbing the signal. Thus, the schema that shortens any redundancies must be read properly by an interpreter to accurately translate the shorthand redundancies back into their longer form. Indeed, it is possible for a signal stream to arrive at a naive receiver in a format that contains sufficient information about how it ought to be decompressed. However, even in this case, the receiver must have prior knowledge about how to decode that signal into its original encoding so as to reveal the redundancies present in the original message that gave it meaning. This prior knowledge means, for example, that one previously knows the chosen encoding-decoding scheme, the programming language, computation model, or the formal theory that explains the relevant physical phenomena which the emitter agent is trying to convey.
In the face of such theoretical challenges, the problem that we tackle in this article is how to encode/decode the message in a compressible signal so as to enable the correct decoding without prior knowledge of the chosen encoding-decoding scheme. We introduce a method based upon the principles of (algorithmic) information theory that is aimed at sweeping over various possible encoding and decoding schemes to test our current limits on signal interpretation as we attempt to reconstruct the original partition (i.e., multidimensional space) in which the original message was given its “meaning” by the emitter agent. This semantic characteristic refers to the (algorithmic) information about the context or real-world correspondents of the emitter that the signal sent by this emitter is trying to convey to the receiver. The real-world correspondents that we particularly investigate are the objects grounded in their respective contexts, that is, embedded into their respective multidimensional spaces as the emitter originally intended.
@@
Zero-knowledge communication occurs when the receiver agent is able to correctly interpret the received signal as the originally intended message sent by the emitter agent, given that the receiver has no knowledge about the encoding-decoding scheme chosen by the emitter. Notice that this condition of no prior information about the emitter agent only applies to encoding-decoding schemes, and consequentially also to the original multidimensional space and object that are unknown to the receiver before any communication takes place. However, as discussed in the Supplementary Information, this zero-knowledge condition does not mean other assumptions regarding the unknown emitter agent are not being considered by the receiver agent. Also, the reader should not confuse zero-knowledge communication (ZKC) with zero-knowledge proof (ZKP), which is commonly studied in cryptography [21, 22, 23]. Actually, one may consider ZKP and ZKC as akin mathematical problems but as diametrically opposing counterparts regarding the acquisition of knowledge (see Sup. Inf.).
@@
As explained in Section 4 and formalised in the Sup. Inf., we estimate the algorithmic probability changes for a message under distinct partitions that result from perturbations. This process then builds a large landscape of possible candidate models (along with their respective approximate complexity values) from which one can infer the best model. We show that the partition (i.e., multidimensional space) for which the message displays the lowest complexity indicates the original partition that the received signal stream encodes. For example in the case of bidimensional spaces, our method consists of finding the bidimensional space $\mathcal{S}_{2}$ for which the algorithmic complexity of the object embedded into $\mathcal{S}_{2}$ (i.e., the algorithmic complexity of the message) is minimised in comparison to the landscape of other complexity values calculated from a sufficiently large number of perturbations on the received signal.
@@
The present work applies AID to the problem of communicating messages by performing a perturbation analysis on the received signal. In this manner, as introduced in the Sup. Inf., our present approach also contributes a universal method that is not reducible to a straightforward application of algorithmic information theory (AIT). As the number of perturbed messages strongly dominates the number of possible best message candidates, our theoretical and empirical results show that perturbation analysis enables the downward spikes displayed in the landscapes of complexities to be more likely to indicate the best message candidates, even though the exact values of the algorithmic complexities are semi-computable (i.e., these values are uncomputable and one can only asymptotically approximate them from above). Although there is this intrinsically observer-dependent“deficiency” in the perturbation analysis that forbids an arbitrarily precise approximation to algorithmic complexity, we demonstrate that this subjective aspect can be eventually cancelled out in the perturbation analysis phase. In fact, it is this theoretical limitation given by semi-computability that, counter-intuitively, enables the perturbation analysis to overcome this and the other subjective aspects (such as an arbitrarily chosen programming language, probability measure, or the computational appliances available) involved in the signal processing and interpretation process—that is, in finding an objective piece of information.
2 Results
2.1 On the perturbation analysis of compressed data
@@
Section A.4 presents empirical results that show our method provides a robust approach to identifying the best way to decode a signal back to its original multidimensional space. Our method consists in sweeping over a large variety of possible dimensional configurations and measuring the resulting information content and algorithmic complexity of the candidate messages by taking the lowest algorithmic complexity configurations under the assumption that the original message is algorithmically not random (i.e., that the original message is compressible). As shown in [29] for graphs, this is because if the original message is random, then the likelihood that a rearrangement (i.e., a new partition) will lead to a low complexity configuration is very low. On the other hand, if the original message is sufficiently compressible (see Section 2.2), and therefore not random, most partitions will lead to configurations of greater algorithmic randomness.
@@
As shown in Figures 11119999 in the Sup. Inf., which apply the method introduced in this article to reconstruct technosignatures, drops (i.e., downward spikes) in the complexity landscapes indicate candidate dimensions for their respective original encoding dimensions, and thus the decoding dimension that would process the signal such that it produces the lowest-complexity message. That is, the original encoding of the message (the one with an interpretable meaning) is the one with the lowest complexity. Our results indicate that sequences that carry meaningful content are those whose complexity quantitatively diverges from the vast majority of the other perturbed messages according to complexity measures. The number of deviations (i.e., perturbations) from the encoded messages that cause the randomness to increase dominates those that do not. In this scenario where a sufficiently large amount of perturbations is taken into account, our results show that an incoming signal with low randomness is more likely to encode a message than one that has more randomness.
@@
In particular, the Block Decomposition Method (BDM) [30] (combining Shannon entropy for long ranges and algorithmic probability for short ranges) was shown to be the complexity measure that was most sensitive and accurate in identifying the original message (image) across the complexity landscapes displayed in Figures $1$, $4$, $6$, $8$, and $9$ in Section A.4. Additionally, as shown in Figure $6$, BDM significantly outperforms Compression and Shannon Entropy in the face of additive noise, whose impact on our message reconstruction method is investigated in Figure $5$. As shown in Figures $2$, $3$, and $7$, algorithmic complexity (as estimated via BDM) remains invariant to low-algorithmic-complexity transformations that preserve the algorithmic content of the data (such as swapping all 0s and 1s or swapping all black and white pixels in an image), as well as some non-linear but algorithmic-information-content-preserving transformations [31, 30].

### v3

# Non-Random Data Encodes Its Geometric and Topological Dimensions

@@ Abstract
Based on the principles of information theory, measure theory, and theoretical computer science, we introduce a signal deconvolution method with a wide range of applications to coding theory, particularly in zero-knowledge one-way communication channels, such as in deciphering messages (i.e., objects embedded into multidimensional spaces) from unknown generating sources about which no prior knowledge is available and to which no return message can be sent. Our multidimensional space reconstruction method from an arbitrary received signal is proven to be agnostic vis-à-vis the encoding-decoding scheme, computation model, programming language, formal theory, the computable (or semi-computable) method of approximation to algorithmic complexity, and any arbitrarily chosen (computable) probability measure. The method derives from the principles of an approach to Artificial General Intelligence (AGI) capable of building a general-purpose model of models independent of any arbitrarily assumed prior probability distribution. We argue that this optimal and universal method of decoding non-random data has applications to signal processing, causal deconvolution, topological and geometric properties encoding, cryptography, and bio- and technosignature detection.
Index Terms:
algorithmic information dynamics, agnostic decoding, randomness, dimension, message structure, meaning, semantics
I Introduction
@@
In this case in which the message is encoded in an uncompressed form, or equivalently in which the message is compressible (therefore non-random), one would have a large standard deviation in the range of information/complexity values for the different ways of decomposing, parsing or perturbing the signal. (See Section III-B and Appendix A). Thus, the schema that shortens any redundancies must be read properly by an interpreter to accurately translate the shorthand redundancies back into their longer form. Indeed, it is possible for a signal stream to arrive at a naive receiver in a format that contains sufficient information about how it ought to be decompressed. However, even in this case, the receiver must have prior knowledge about how to decode that signal into its original encoding so as to reveal the redundancies present in the original message that gave it meaning. This prior knowledge means, for example, that one previously knows the chosen encoding-decoding scheme, the programming language, computation model, or the formal theory that explains the relevant physical phenomena which the emitter agent is trying to convey.
In the face of such theoretical challenges, the problem that we tackle in this article is how to encode/decode the message in a compressible signal so as to enable the correct decoding without prior knowledge of the chosen encoding-decoding scheme. We introduce a method based upon the principles of (algorithmic) information theory that is aimed at sweeping over various possible encoding and decoding schemes to test our current limits on signal interpretation as we attempt to reconstruct the original partition (or in general, the multidimensional space or structure) in which the original message was given its “meaning” by the emitter agent. This semantic characteristic refers to the (algorithmic) information about the context or real-world correspondents of the emitter that the signal sent by this emitter is trying to convey to the receiver. The real-world correspondents that we particularly investigate are the objects grounded in their respective contexts, that is, embedded into their respective multidimensional spaces (or structures) as the emitter originally intended.
@@
Zero-knowledge communication occurs when the receiver agent is able to correctly interpret the received signal as the originally intended message sent by the emitter agent, given that the receiver has no knowledge about the encoding-decoding scheme chosen by the emitter. Notice that this condition of no prior information about the emitter agent only applies to encoding-decoding schemes—in addition to other language, computational, and mathematical modeling characteristics, such as the programming language, computational model, and probability measure—, and (consequentially) also to the original multidimensional space and object that are unknown to the receiver before any communication takes place. However, as discussed in the Supplementary Material provided along with the present article, this zero-knowledge condition does not mean other assumptions regarding the unknown emitter agent are not being considered by the receiver agent. Also, the reader should not confuse zero-knowledge communication (ZKC) with zero-knowledge proof (ZKP), which is commonly studied in cryptography [21, 22, 23]. Actually, one may consider ZKP and ZKC as akin mathematical problems but as diametrically opposing counterparts regarding the acquisition of knowledge (see also Appendix D).
@@
As explained in Section III and formalised in the Sup. Mat., we estimate the algorithmic probability changes for a message under distinct partitions that result from perturbations. This process then builds a large landscape of possible candidate models (along with their respective approximate complexity values) from which one can infer the best model. We show that the partition (i.e., multidimensional space) for which the message displays the lowest complexity indicates the original partition that the received signal stream encodes. For example in the case of bidimensional spaces, our method consists of finding the bidimensional space $\mathcal{S}_{2}$ for which the algorithmic complexity of the object embedded into $\mathcal{S}_{2}$ (i.e., the algorithmic complexity of the message) is minimised in comparison to the landscape of other complexity values calculated from a sufficiently large number of perturbations on the received signal.
@@
In this manner, our present approach also contributes a universal method that is not reducible to a straightforward application of algorithmic information theory (AIT). As the number of perturbed messages strongly dominates the number of possible best message candidates, our theoretical and empirical results show that perturbation analysis enables the downward spikes displayed in the landscapes of complexities to be more likely to indicate the best message candidates, even though the exact values of the algorithmic complexities are semi-computable (i.e., these values are uncomputable and one can only asymptotically approximate them from above). Although there is this intrinsically observer-dependent “deficiency” in the perturbation analysis that forbids an arbitrarily precise approximation to algorithmic complexity, we demonstrate that this subjective aspect can be eventually cancelled out in the perturbation analysis phase. In fact, it is this theoretical limitation given by semi-computability that, counter-intuitively, enables the perturbation analysis to overcome this and the other subjective aspects (such as an arbitrarily chosen programming language, probability measure, or the computational appliances available) involved in the signal processing and interpretation process—that is, in finding an objective piece of information.
II Empirical Results
II-A On the perturbation analysis of compressed data
@@
Our (structural) perturbation method consists in sweeping over a large variety of possible dimensional configurations and measuring the resulting information content and algorithmic complexity of the candidate messages by taking the lowest algorithmic complexity configurations under the assumption that the original message is algorithmically not random (i.e., that the original message is compressible). As shown in [29] for graphs, this is because if the original message is random, then the likelihood that a rearrangement (i.e., a new partition) will lead to a low complexity configuration is very low. On the other hand, if the original message is sufficiently compressible (see Sections III-B and III-C), and therefore not random, most partitions will lead to configurations of greater algorithmic randomness.
@@
As shown in [30] which applies the method discussed in this article to reconstruct technosignatures, drops (i.e., downward spikes) in the complexity landscapes indicate candidate dimensions for their respective original encoding dimensions, and thus the decoding dimension that would process the signal such that it produces the lowest-complexity message. That is, the original encoding of the message (the one with an interpretable meaning) is the one with the lowest complexity. Our results indicate that sequences that carry meaningful content are those whose complexity quantitatively diverges from the vast majority of the other perturbed messages according to complexity measures. The number of deviations (i.e., perturbations) from the encoded messages that cause the randomness to increase dominates those that do not. In this scenario where a sufficiently large amount of perturbations is taken into account, our results show that an incoming signal with low randomness is more likely to encode a message than one that has more randomness.
@@
In particular, the Block Decomposition Method (BDM) [31] (combining Shannon entropy for long ranges and algorithmic probability for short ranges) was shown in [30] to be the complexity measure that was most sensitive and accurate in identifying the original message (image) across the complexity landscapes.
Additionally, BDM significantly outperforms Compression and Shannon Entropy in the face of additive noise. As shown in [30], algorithmic complexity (as estimated via BDM) remains invariant to low-algorithmic-complexity transformations that preserve the algorithmic content of the data (such as swapping all 0s and 1s or swapping all black and white pixels in an image), as well as some non-linear but algorithmic-information-content-preserving transformations [32, 31].